HUBBLE SPACE TELESCOPE DAILY REPORT #1102 PERIOD COVERED: 0700 EST Daily Status Report as of 083/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1 5277 (GHRS Side One Spectral Lamp Test: Cycle 4) The High Resolution Spectrograph (Side One) was used to make internal spectral lamp wavelength calibrations. The data analysis will provide dispersion constants, carrousel and Y-deflection functions, and spectral resolving power for Side 1. The GHRS internals appeared to complete without apparent anomaly, except for a carrousel lock failure. Six observations were received and displayed. They all were reported to be satisfactory. 1.2 Completed WF/PC-2 5476 (Continuum Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to attempt snapshot images of a radio galaxy as part of a continuum survey and to study the galaxy's morphology. The baseline guide star acquisition failed. The observations only received a shutter toggle instead of the planned 300 second exposures. 1.3 Completed GHRS/2 5596 (GHRS Side 2 Sensitivity Calibration (1ST Order Gratings): Cycle 4) The High Resolution Spectrograph (Side Two) was used to measure its sensitivity to provide information on temporal variations in throughput since the installation of COSTAR. Observations were made of the standard test object MU-COLUMBIA. All of the guide star acquisitions achieved fine lock. The onboard target acquisition appeared to work as desired. No anomalies were noted. 1.4 Completed GHRS/2 5164 (Coronal Line Survey in Late Type Stars: Cycle 4 High Priority) The High Resolution Spectrograph (Side Two) was used to measure coronal lines in the late-type star HD22468 (aka, V711 Tau) which is an RS CVn-type variable star. The onboard target acquisitions were monitored in the OSS, and they appeared to be successful. Some observations still remain on the science tape recorder, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 There was a Loss of Lock (LOL) at 082/1354Z. FGS #3 failed to acquire the designated guide star. The probable cause was a roll update following a Type 2 slew, instead of a full axis update. The subsequent acquisition was successful. Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled :12 Successful :12 2.3 Significant HSTARs: None 2.4 Operations Notes: Another anomaly-free day was experienced, as every HST subsystem reported no operational problems. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of GHRS/2 5596 (GHRS Side 2 Sensitivity Calibration (1st Order Gratings): Cycle 4) 3.2 Complete WF/PC-2 4709 (WF/PC-2 Internal Monitor) 3.3 Complete WF/PC-2 5176 (Lyman-Alpha Region of QSOs With Strong Absorption Lines: Cycle 4) 3.4 Complete FOC/96 5517 (FOC Absolute Sensitivity: Cycle 4) 3.5 Complete Two Set of WF/PC-2 5370 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) 3.6 Complete WF/PC-2 5091 (GTO High Latitude Parallel Program: Cycle 4) 3.7 Complete GHRS/1 5592 (GHRS ECH-A Sensitivity Calibration: Cycle 4) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-